# SAEmnesia: Erasing Concepts in Diffusion Models with Sparse Autoencoders

### v1

@@ Abstract
Effective concept unlearning in text-to-image diffusion models requires precise localization of concept representations within the model’s latent space. While sparse autoencoders successfully reduce neuron polysemanticity (i.e., multiple concepts per neuron) compared to the original network, individual concept representations can still be distributed across multiple latent features, requiring extensive search procedures for concept unlearning. We introduce SAEmnesia, a supervised sparse autoencoder training method that promotes one-to-one concept-neuron mappings through systematic concept labeling, mitigating feature splitting and promoting feature centralization. Our approach learns specialized neurons with significantly stronger concept associations compared to unsupervised baselines. The only computational overhead introduced by SAEmnesia is limited to cross-entropy computation during training. At inference time, this interpretable representation reduces hyperparameter search by 96.67% with respect to current approaches. On the UnlearnCanvas benchmark, SAEmnesia achieves an 9.22% improvement over the state-of-the-art. In sequential unlearning tasks, we demonstrate superior scalability with a 28.4% improvement in unlearning accuracy for 9-object removal.
@@ 1 Introduction
Text-to-image diffusion models have achieved remarkable success in generating high-quality images from textual descriptions, with applications across diverse domains [33]. However, they can also produce harmful, inappropriate, or copyrighted content when given specific prompts, raising serious safety concerns. This has spurred growing interest in machine unlearning methods, which aim to selectively remove undesired concepts from trained models while preserving their generative abilities [43].
@@
Figure 1: SAEmnesia enables precise concept-level manipulation: each latent activates for a single concept (monosemanticity), and each concept is embedded in a single latent (feature centralization). So, to erase a target concept, we only need to steer a single latent. The removed concepts correctly disappear in the diagonal images (Architectures, Bears, Flames), while remaining present in the non-diagonal ones, thereby preserving the fidelity and diversity of unrelated content.
@@
A core challenge in concept unlearning is identifying where and how concepts are represented inside these models, especially when the number of trainable parameters is very high. Additionally, a single neuron can be active for many different unrelated concepts. This phenomenon is known as polysemanticity, making interpretability even more challenging. Mechanistic interpretability (MI) seek to understand the internal workings of neural networks by analyzing their components and representations. Sparse Autoencoders (SAEs) provide a particularly effective MI tool by decomposing model activations into sparse and interpretable concept-level features [5]. Note that, in this work, we adopt the notion of features as the fundamental units of neural network representations that cannot be further decomposed into simpler independent factors, as defined by  Bereska and Gavves [4]. Neural networks can capture natural abstractions [7] through their learned features, which serve as building blocks of their internal representations, aiming to capture the concepts underlying the data. For simplicity, we use the terms “concepts” and “features” interchangeably, as well as “neurons” and “latents”. SAEs aim to learn monosemantic latents, meaning that they activate almost exclusively for a specific concept. On the other hand, to improve models’ interpretability even further, individual concepts should not be spread across many latents, aiming for a one-to-one mapping. Yet, in practice, multiple latents often respond to the same concept, this phenomenon is also known as feature splitting [5]. This distributed nature poses two key challenges: (i) it necessitates exhaustive searches across many latent combinations to find the right subset to modify, making unlearning computationally expensive; and (ii) overlapping latents blur concept boundaries, so interventions risk unintended side effects on related concepts.
@@
To overcome feature splitting, we introduce SAEmnesia, that enriches the SAE training framework enforcing one-to-one mappings between concepts and latents through supervised labeling. Therefore, our method achieves feature centralization, localizing each concept into a single latent feature and thereby preventing splitting across multiple latents. This binding simplifies mechanistic unlearning, thanks to a precise single-latent intervention. Figure 1 illustrates the effect of SAEmnesia for concept removal in generative models, demonstrating the possibility of precise control over what the model generates, while retaining overall quality and diversity. On the UnlearnCanvas benchmark [43], SAEmnesia achieves an 9.22% improvement over the state-of-the-art mechanistic approaches [9]. In sequential unlearning tasks, we demonstrate superior scalability with a 28.4% improvement in unlearning accuracy for 9-object removal. Furthermore, at inference time, this interpretable representation reduces hyperparameter search by 96.67%.
@@ 2 Related Work
Machine Unlearning in Diffusion Models. The term and problem statement for machine unlearning was first introduced by Cao and Yang [6], who transformed neural network models through additional simple layers into formats where output is a summation of independent features, allowing unlearning by blocking selected summation weights or nodes. However, recent works focusing on unlearning for diffusion models typically employ fine-tuning approaches to unlearn specific concepts.
@@
EDiff [40] formulates this problem as bi-level optimization, while ESD [14] leverages negative classifier-free guidance for concept removal. FMN [41] introduces a re-steering loss applied only to attention layers, and SalUn [11] and SHS [39] select parameters to adapt through saliency maps or connection sensitivity. SA [19] replaces unwanted data distribution with surrogate distributions, with an extension to selected anchor concepts in CA [24]. SPM [28] takes a different approach, using small linear adapters added after each linear and convolutional layer to directly block unwanted content propagation. In contrast to these fine-tuning approaches, our work leverages SAEs to achieve unlearning through interpretable feature manipulation during inference, without modifying the base model weights and providing full transparency into which specific features are being targeted for removal.
By contrast, methods that do not rely on fine-tuning include SEOT [26], which eliminates unwanted content from text embeddings, and UCE [15], which adapts cross-attention weights using closed-form solutions. Unlike these approaches that modify prompt embeddings or alter model weights, our method operates through interpretable feature interventions on internal activations during inference.
@@
Beyond individual methods, recent surveys synthesize objectives, taxonomies, and evaluation protocols for generative model unlearning, offering broader context for method design and assessment [12, 8]. Furthermore, recent analyses highlight instability and concept resurgence after unlearning (such as revival under subsequent fine-tuning or adversarial prompting) reinforcing the need for interpretable and stable interventions [16, 27, 34].
@@
SAE Background. Our objective is to enable effective concept unlearning in diffusion models by selectively removing unwanted concepts while preserving generative quality. To achieve this, we decompose the high-dimensional, entangled activations from Stable Diffusion (SD) into sparse, interpretable feature directions that correspond to meaningful visual concepts. SAEs serve as the key tool for this decomposition, enabling us to map individual neurons to specific semantic concepts and subsequently intervene on them for targeted unlearning. A standard single-layer ReLU SAE [30] operates on $d$-dimensional activation vectors. Let $\mathbf{x}\in\mathbb{R}^{d}$ denote the input activation vector and $n$ be the latent dimension, typically set to $d$ multiplied by a positive expansion factor. The encoder and decoder are defined as [5]:
$\begin{split}\mathbf{v}&=\text{ReLU}(W_{enc}(\mathbf{x}-\mathbf{b}_{pre})+\mathbf{b}_{enc})\\ \hat{\mathbf{x}}&=W_{dec}\mathbf{v}+\mathbf{b}_{pre},\end{split}$ (1)
@@
where $W_{enc}\in\mathbb{R}^{n\times d}$ and $W_{dec}\in\mathbb{R}^{d\times n}$ are the encoder and decoder weight matrices respectively, and $\mathbf{b}_{pre}\in\mathbb{R}^{d}$ and $\mathbf{b}_{enc}\in\mathbb{R}^{n}$ are learnable bias terms.
@@
TopK SAEs. In our work, we employ TopK SAEs [29] that provide enhanced sparsity control. The TopK activation function identifies the $k$ largest pre-activation values and sets all others to zero, ensuring sparsity while preserving the most significant features. Following recent advances, we adopt this approach where the encoder becomes:
$\begin{split}\mathbf{z}&=\text{TopK}(\mathbf{v}),\\ \hat{\mathbf{x}}&=W_{dec}\mathbf{z}+\mathbf{b}_{pre},\end{split}$ (2)
while the decoder remains unchanged. Here, $\mathbf{v}$ represents the pre-TopK activations, while $\mathbf{z}$ represents the post-TopK sparse activations. This modification retains only the $k$ largest latent activations for each vector $\mathbf{x}$, providing better control over sparsity compared to standard ReLU activations.
@@
Training Objective. The TopK SAE loss combines reconstruction error with an auxiliary loss to prevent dead latents:
[…]
@@
where $\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2}$ is the reconstruction error and $\mathcal{L}_{aux}$ is an auxiliary loss using only the largest $k_{aux}$ feature activations that have not fired on a large number of training samples (so-called dead latents). The auxiliary loss prevents dead latents from occurring and is scaled by coefficient $\alpha$.
@@
Interpretability of Vision Models. SAEs have recently gained traction as a tool for uncovering human-interpretable structure in high-dimensional representations. Early applications focused on discriminative models, such as interpreting features in CLIP [13, 10, 32] or traditional classifiers [36, 17]. More recent work has extended SAEs to vision–language settings, enabling tasks such as hallucination mitigation [21] and interpretable report generation [1]. Across these domains, SAEs provide explicit concept–neuron mappings that move beyond post-hoc explanations and allow more precise control over learned representations. Unlike previous work on discriminative models, our approach uses SAEs with diffusion models to directly identify and control specific concepts.
@@
Within diffusion models, interpretability research has primarily examined how semantic information propagates through the architecture. Studies have identified meaningful directions in UNet bottlenecks [25, 31, 18, 20], analyzed cross-attention to link prompts with spatial activations [37, 2, 3], and even used intermediate text encoder states for generation [38]. More recent work has highlighted interpretable feature detectors in SDXL-Turbo [35] and timestep-specific visual features in unconditional sampling [23]. While these advances shed light on model internals, they generally stop short of enabling targeted interventions. Our approach yields sparse representations that support both interpretation and controllable unlearning of specific concepts.
@@
SAE-Based Unlearning. Recent work has explored applying SAEs to diffusion models for concept manipulation. These approaches differ from ours in their intervention points and training strategies. Kim and Ghadiyaram [22] introduced Concept Steerers, training SAEs on text embedding representations to identify concept-specific directions before cross-attention processing. While their approach achieves effective concept manipulation, working only on text encoders can lead to suboptimal results, especially when facing adversarial attacks that exploit deeper model representations. Cywiński and Deja [9] introduced SAeUron, a post-cross-attention approach using SAEs trained on diffusion model activations in an unsupervised manner. While SAeUron achieves state-of-the-art performance on UnlearnCanvas, its unsupervised training creates weak concept-latent associations. This means that concepts are still represented across multiple neurons. This distributed representation requires computationally expensive feature threshold searches to identify which combination of latent features must be modified to unlearn each concept.
Our approach addresses these limitations through supervised training that enforces strong one-to-one concept-latent relationships, enabling precise single-neuron interventions.
@@ 3 Methodology
Figure 2: SAEmnesia comprises two phases: (i) establishing sparse representations via standard unsupervised SAE training, (ii) applying supervised losses to strengthen specific concept-neuron associations.
Unlike unsupervised SAE training methods [5] that require post-hoc discovery of concept-relevant features, our supervised approach directly enforces concept-latent assignments during training to achieve stronger one-to-one mappings. As shown in Figure 2, SAEmnesia consists of two phases: (i) establishing sparse representations through standard unsupervised SAE training, (ii) applying supervised losses to strengthen specific concept-neuron associations.
SAEmnesia for Diffusion Models. We apply SAEmnesia to diffusion models by training on activations extracted from every timestep $t$ of the denoising diffusion process. These activations are obtained from the cross-attention blocks of the diffusion model and form feature maps. Each feature map extracted at timestep $t$ is a spatially structured tensor of shape $\mathbf{F}_{t}\in\mathbb{R}^{h\times w\times d}$, where $h$ and $w$ denote the height and width of the feature map, and $d$ is the dimensionality of each feature vector. Each spatial position within the feature map corresponds to a patch in the input image. As a single SAE training sample, we consider an individual $d$-dimensional feature vector $\mathbf{x}$, disregarding the information about its spatial position. Therefore, from each feature map, we obtain $h\times w$ training samples.
Concept-Latent Assignment. This assignment is required in the second phase, in which we introduce supervised training that directly assigns concepts to specific latents. To determine which latent should be assigned to each concept during the supervised phase and to validate the quality of these assignments after training, we utilize the score function [9], defined in Equation 4, measuring feature-concept correspondence. Given a dataset of activations $D=D_{c}\cup D_{\neg c}$, where $D_{c}$ contains data of the target concept $c$ and $D_{\neg c}$ does not, the score function is defined as:
@@
Therefore, for each concept $c$, we assign it to the latent index $i_{c}$ with the highest score, creating a mapping $\Phi:\mathcal{C}\rightarrow\{1,2,...,n\}$, where $\mathcal{C}$ is the set of concepts to unlearn. For training samples containing multiple concepts, we assign multiple target indices corresponding to each present concept.
@@
$\mathcal{L}_{\text{{SAEmnesia}}}=\mathcal{L}_{unsupSAE}+\beta\mathcal{L}_{supSAE}+\lambda\mathcal{L}_{L1}.$ (6)
@@
Where $\mathcal{L}_{unsupSAE}$ is the loss as defined in Equation (3), $\mathcal{L}_{L1}$ is the sparsity regularization term, and $\mathcal{L}_{supSAE}$ is the supervised loss that enforces the desired concept-latent relationships. $\mathcal{L}_{supSAE}$ consist of our Concept Assignment (CA) loss with a weighted additional Orthogonality Constraint (OC) loss:
$\mathcal{L}_{supSAE}=\mathcal{L}_{CA}+\gamma\mathcal{L}_{OC}.$ (7)
@@
Concept Assignment Loss. Unlike traditional approaches that only apply reconstruction loss globally across all latents, our method requires additional supervision to strengthen concept-latent bonds that enable one-to-one mappings. We employ the CA loss that is computed exclusively for latents that are assigned to concepts present in each training sample, encouraging these specific latents to activate strongly.
For each training sample $i$, we identify the set of target positions $\mathcal{T}=\{j:$ latent $j$ is assigned to a concept present in sample i}i\}. The CA loss measures how much these assigned latents activate when their corresponding concepts are present:
$\begin{split}\mathcal{L}_{\text{CA}}&=\frac{1}{|\mathcal{T}|}\sum_{(i,j)\in\mathcal{T}}\Big[-\log(\sigma(v^{i}_{j}))\Big],\end{split}$ (8)
where $v^{i}_{j}$ represents the pre-activation (logit) for latent $j$ in sample $i$, and $\sigma(\cdot)$ is the sigmoid function. This loss directly penalizes assigned latents that fail to achieve strong activation when their concepts are present, while completely ignoring latents not assigned to concepts in each sample.
Orthogonality Constraint. A critical challenge in concept unlearning is ensuring that removing one concept (e.g., a specific object) does not inadvertently affect related but distinct concepts (e.g., artistic styles). This problem arises when latents representing different concepts become correlated, leading to entangled representations where modifications to one concept cascade into unintended changes in others. For instance, if a latent assigned to ”cars” becomes correlated with a latent representing ”impressionist style,” then unlearning cars might also diminish the model’s ability to generate impressionist artwork.
To address this entanglement issue, we introduce an across-batch orthogonality constraint that explicitly decorrelates the activation patterns of object and style latents:
$\mathcal{L}_{OC}=\frac{1}{|\mathcal{O}|\cdot|\mathcal{S}|}\sum_{o\in\mathcal{O}}\sum_{s\in\mathcal{S}}\rho^{2}(a_{o},a_{s}),$ (9)
where $\mathcal{O}$ and $\mathcal{S}$ represent the sets of assigned object and style latent indices, $a_{o}=[v^{1}_{o},v^{2}_{o},\ldots,v^{B}_{o}]^{T}$ denotes the pre-TopK activation vector across the batch for latent labeled with object $o$ (similarly for $a_{s}$), and $\rho(a_{o},a_{s})$ is the Pearson correlation coefficient. This constraint minimizes correlations between activation patterns of different latents across the batch, ensuring the information of different concepts remains separated.
Sparsity Regularization. To encourage sparse activations in the latent representation, we incorporate an L1 regularization term that penalizes the magnitude of latent activations:
$\mathcal{L}_{L1}=\frac{1}{|N|}\sum_{j=1}^{|N|}\|v_{j}\|_{1},$ (10)
@@
where $|N|$ is the number of latents and $v_{j}$ represents the activation for latent $j$. This sparsity constraint promotes the emergence of interpretable features by encouraging most latents to remain inactive for any given input, thereby improving the disentanglement of learned representations.

### v2

# SAEmnesia: Erasing Concepts in Diffusion Models with Supervised Sparse Autoencoders

@@ Abstract
Concept unlearning in diffusion models is hampered by feature splitting, where concepts are distributed across many latent features, making their removal challenging and computationally expensive. We introduce SAEmnesia, a supervised sparse autoencoder framework that overcomes this by enforcing one-to-one concept-neuron mappings. By systematically labeling concepts during training, our method achieves feature centralization, binding each concept to a single, interpretable neuron. This enables highly targeted and efficient concept erasure. SAEmnesia reduces hyperparameter search by 96.7% and achieves a 9.2% improvement over the state-of-the-art on the UnlearnCanvas benchmark. Our method also demonstrates superior scalability in sequential unlearning, improving accuracy by 28.4% when removing nine objects, establishing a new standard for precise and controllable concept erasure. Moreover, SAEmnesia mitigates the possibility of generating unwanted content under adversarial attack and effectively removes nudity when evaluated with I2P.
@@ 1 Introduction
Text-to-image diffusion models have achieved remarkable success in generating high-quality images from textual descriptions, with applications across diverse domains [rombach2022high]. However, they can also produce harmful, inappropriate, or copyrighted content when given specific prompts, raising safety concerns. This has spurred growing interest in machine unlearning, which aims to selectively remove undesired concepts from trained models while preserving their generative abilities [zhang2024unlearncanvas].
@@
Figure 1: SAEmnesia enables precise concept-level manipulation: each latent activates for a single concept (monosemanticity), and each concept is embedded in a single latent (feature centralization). So, to erase a target concept, we only need to steer a single latent. The removed concepts correctly disappear in the diagonal images (“Architectures”, “Bears”, “Sandwiches”) while the corresponding style is preserved. Note that they remain present in the non-diagonal ones, thereby preserving the fidelity and diversity when unlearning unrelated content.
@@
A core challenge in concept unlearning is identifying where and how concepts are represented inside these models. Each neuron can encode multiple unrelated concepts simultaneously. This phenomenon is known as polysemanticity, making interpretability even more challenging. Mechanistic interpretability (MI) seeks to understand the internal workings of neural networks by analyzing their representations. Sparse Autoencoders (SAEs) provide a particularly effective MI tool by decomposing model activations into sparse and interpretable concept-level features [bricken2023monosemanticity]. In this work, we adopt the notion of features as the fundamental units of neural network representations that cannot be further decomposed into simpler independent factors, as defined by  bereska2024mechanistic. Neural networks can capture natural abstractions [Chan2023NaturalAbstractions] through their learned features, which serve as building blocks of their internal representations, aiming to capture the concepts underlying the data. For simplicity, we use the terms “concepts” and “features” interchangeably, as well as “neurons” and “latents”. SAEs aim to learn monosemantic latents, meaning that they activate almost exclusively for a specific concept. On the other hand, to improve models’ interpretability even further, individual concepts should not be spread across many latents, aiming for a one-to-one mapping. Yet, in practice, multiple latents often respond to the same concept, this phenomenon is also known as feature splitting [bricken2023monosemanticity]. This means modifying one concept requires changing multiple neurons. This distributed nature poses two key challenges: (i) exhaustive searches across many latent combinations to find the right subset to modify, making unlearning computationally expensive; and (ii) overlapping latents blur concept boundaries, so interventions risk unintended side effects on related concepts.
@@
To overcome feature splitting, we introduce SAEmnesia, that enriches the SAE training framework enforcing one-to-one mappings between concepts and latents through supervised labeling. Therefore, our method achieves feature centralization, localizing each concept into a single latent and thereby preventing splitting across multiple neurons. This binding simplifies mechanistic unlearning, thanks to a precise single-latent intervention. Fig. 1 illustrates the effect of SAEmnesia for concept removal in generative models, demonstrating the possibility of precise control over what the model generates, while retaining overall quality and diversity. On the UnlearnCanvas benchmark [zhang2024unlearncanvas], SAEmnesia achieves an 9.22% improvement over the state-of-the-art mechanistic approaches [cywinski2025saeuron]. In sequential unlearning tasks, we demonstrate superior scalability with a 28.4% improvement in unlearning accuracy for 9-object removal. Furthermore, at inference, this interpretable representation reduces hyperparameter search by 96.67%.
In summary, our key contributions are as follows: (i) we introduce a supervised sparse autoencoder that explicitly enforces one-to-one concept–latent mappings, eliminating feature splitting and enabling transparent, interpretable control over concept representations; (ii) we show that this structure makes concept erasure significantly more efficient: each concept can be removed by steering a single latent, substantially reducing inference-time hyperparameter search; (iii) we achieve state-of-the-art performance on the UnlearnCanvas benchmark and demonstrate improved sequential unlearning, stronger robustness to adversarial attacks, and more effective NSFW-content suppression.
@@ 2 Related Work
Machine unlearning in diffusion models. Machine unlearning was first introduced by cao2015towards, who transformed neural network models through additional simple layers into formats where output is a summation of independent features, allowing unlearning by blocking selected summation weights or nodes. However, recent works focusing on unlearning for diffusion models typically employ fine-tuning approaches to unlearn specific concepts. EDiff [wu2024erasediff] formulates this problem as bi-level optimization, while ESD [gandikota2023erasing] leverages negative classifier-free guidance for concept removal. FMN [zhang2024forget] introduces a re-steering loss applied only to attention layers, and SalUn [fan2023salun] and SHS [wu2024scissorhands] select parameters to adapt through saliency maps or connection sensitivity. SA [heng2023selective] replaces unwanted data distribution with surrogate distributions, with an extension to selected anchor concepts in CA [kumari2023ablating]. SPM [lyu2024one] takes a different approach, using small linear adapters added after each linear and convolutional layer to directly block unwanted content propagation. Methods that do not rely on fine-tuning include SEOT [li2024get], which removes unwanted content from text embeddings, and UCE [gandikota2024unified], which adapts cross-attention weights using closed-form solutions. In contrast to these approaches, our work leverages SAEs to achieve unlearning through interpretable feature manipulation during inference, without modifying the base model weights and providing full transparency into which specific features are being targeted for removal.
@@
Beyond individual methods, recent surveys synthesize objectives, taxonomies, and evaluation protocols for generative model unlearning, offering broader context for method design and assessment [feng2025surveygenmu, chen2025unlearning_survey]. Furthermore, recent analyses highlight instability and concept resurgence after unlearning (such as revival under subsequent fine-tuning or adversarial prompting) reinforcing the need for interpretable and stable interventions [george2025illusion, lu2025when, suriyakumar2024resurgence].
@@
SAEs background. Our aim is to enable effective concept unlearning in diffusion models by selectively removing unwanted concepts while preserving generative quality. To achieve this, we decompose the high-dimensional, entangled activations from Stable Diffusion (SD) into sparse, interpretable feature directions that correspond to meaningful visual concepts. SAEs serve as the key tool for this decomposition, enabling us to map individual neurons to specific semantic concepts and subsequently intervene on them for targeted unlearning. A standard single-layer ReLU SAE [Olshausen1997SparseCW] operates on $d$-dimensional activation vectors. Let $\mathbf{x}\in\mathbb{R}^{d}$ denote the input activation vector and $n$ be the latent dimension, typically set to $d$ multiplied by a positive expansion factor. The encoder and decoder are defined as [bricken2023monosemanticity]:
$\begin{split}\mathbf{v}&=\text{ReLU}(\mathbf{W}_{\text{enc}}(\mathbf{x}-\mathbf{b}_{\text{pre}})+\mathbf{b}_{\text{enc}})\\ \hat{\mathbf{x}}&=\mathbf{W}_{\text{dec}}\mathbf{v}+\mathbf{b}_{\text{pre}},\end{split}$ (1)
@@
where $\mathbf{v}$ is the sparse hidden representation, $\hat{\mathbf{x}}$ is the reconstructed input, $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{n\times d}$ and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times n}$ are the encoder and decoder weight matrices respectively, and $\mathbf{b}_{\text{pre}}\in\mathbb{R}^{d}$ and $\mathbf{b}_{\text{enc}}\in\mathbb{R}^{n}$ are learnable bias terms.
@@
TopK SAEs. In our work, we employ TopK SAEs [Makhzani2013kSparseA] that provide enhanced sparsity control. The TopK activation function identifies the $k$ largest pre-activations and sets all others to zero, ensuring sparsity while preserving the most significant features:
$\begin{split}\mathbf{z}&=\text{TopK}(\mathbf{v}),\\ \hat{\mathbf{x}}&=\mathbf{W}_{\text{dec}}\mathbf{z}+\mathbf{b}_{\text{pre}}.\end{split}$ (2)
Here, $\mathbf{v}$ represents the pre-TopK activations, while $\mathbf{z}$ represents the post-TopK sparse activations.
@@
Training objective. Given a mini-batch of size $B$, the TopK SAE loss combines reconstruction error with an auxiliary loss to prevent dead latents:
$\mathcal{L}_{\text{unsupSAE}}=\frac{1}{B}\sum_{b=1}^{B}\big\|\mathbf{x}^{(b)}-\hat{\mathbf{x}}^{(b)}\big\|_{2}^{2}+\alpha\,\mathcal{L}_{\text{aux}},$ (3)
@@
where $\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2}$ is the reconstruction error and $\mathcal{L}_{\text{aux}}$ is an auxiliary loss using only the largest $k_{\text{aux}}$ feature activations that have not fired on a large number of training samples (so-called dead latents). The auxiliary loss prevents dead latents from occurring and is scaled by coefficient $\alpha$.
@@
Interpretability of vision models. SAEs have recently gained traction as a tool for uncovering human-interpretable structure in high-dimensional representations. Early applications focused on discriminative models, such as interpreting features in CLIP [fry2024towards, daujotas2024case, rao2024discover] or traditional classifiers [szegedy2015going, gorton2024missing]. More recent work has extended SAEs to vision–language settings, enabling tasks such as hallucination mitigation [jiang2024interpreting] and interpretable report generation [abdulaal2024x]. Across these domains, SAEs provide explicit concept–neuron mappings that move beyond post-hoc explanations and allow more precise control over learned representations. Unlike previous work on discriminative models, our approach uses SAEs with diffusion models to directly identify and control specific concepts.
@@
Within diffusion models, interpretability research has primarily examined how semantic information propagates through the architecture. Studies have identified meaningful directions in UNet bottlenecks [kwon2022diffusion, park2023understanding, hahm2024isometric, ijishakin2024h, Kim_2025_ICCV], analyzed cross-attention to link prompts with spatial activations [tang2022daam, basu2023localizing, basu2024mechanistic], and even used intermediate text encoder states for generation [toker2024diffusion]. Although these advances shed light on the internals of the model, they generally do not enable targeted interventions. Our approach yields sparse representations that support both interpretation and controllable unlearning of specific concepts.
@@
SAE-Based unlearning. Recent work has explored applying SAEs to diffusion models for concept manipulation. kim2025concept introduced Concept Steerers, training SAEs on text embedding representations to identify concept-specific directions before cross-attention processing. While their approach achieves effective concept manipulation, working only on text encoders can lead to suboptimal results, especially when facing adversarial attacks that exploit deeper model representations. cywinski2025saeuron introduced SAeUron, a post-cross-attention approach using SAEs trained on diffusion model activations in an unsupervised manner. While SAeUron achieves state-of-the-art performance on UnlearnCanvas, its unsupervised training creates weak concept-latent associations. This means that concepts are still represented across multiple neurons. This distributed representation requires computationally expensive feature threshold searches to identify which combination of latent features must be modified to unlearn each concept. In contrast, we only need to steer a single feature per concept.
@@ 3 Methodology
Figure 2: SAEmnesia pipeline. Training comprises two phases: (i) establishing sparse representations via standard unsupervised SAE training, (ii) applying supervised losses to strengthen specific concept-neuron associations. During inference, we need to steer a single latent per concept.
Unlike unsupervised SAE training methods [bricken2023monosemanticity] that require post-hoc discovery of concept-relevant features, our supervised approach directly enforces concept-latent assignments during training to achieve stronger one-to-one mappings (see Fig. 2). Although our approach requires supervision, the labels come at no additional cost, as they are directly derived from the text prompts used to generate or condition the images: i.e., the same concepts that the SAE aims to forget.
SAEmnesia for diffusion models. We apply SAEmnesia to diffusion models by training on activations extracted from every timestep $t$ of the denoising diffusion process. These activations are obtained from the cross-attention blocks of the diffusion model and form feature maps. Each feature map extracted at timestep $t$ is a spatially structured tensor of shape $\mathbf{F}_{t}\in\mathbb{R}^{h\times w\times d}$, where $h$ and $w$ denote the height and width of the feature map, and $d$ is the dimensionality of each feature vector. Each spatial position within the feature map corresponds to a patch in the input image. As a single SAE training sample, we consider an individual $d$-dimensional feature vector $\mathbf{x}$, disregarding the information about its spatial position. Therefore, from each feature map, we obtain $h\times w$ training samples. Note that our method is architecture agnostic, meaning that it can be transferred to various text-to-image (T2I) models.
Concept-latent assignment. Here we introduce supervised training that directly assigns concepts to specific latents. To determine which latent should be assigned to each concept during the supervised phase and to validate the quality of these assignments after training, we utilize the score function [cywinski2025saeuron], defined in Eq. 4, measuring feature-concept correspondence. Given a dataset of activations $D=D_{c}\cup D_{\neg c}$, where $D_{c}$ contains data of the target concept $c$ and $D_{\neg c}$ does not, the score function is defined as:
@@
Given a set $\mathcal{C}=\{c_{1},\ldots,c_{K}\}$ of concepts to unlearn, we define a mapping $\Phi:\mathcal{C}\to\{1,\ldots,n\}$ assigning each concept $c$ to the latent index $\Phi(c)=i_{c}$ with the highest score. For training samples containing multiple concepts, we assign multiple target indices corresponding to each present concept. By enforcing one-to-one concept-to-latent mappings, the cardinality of the SAE’s hidden layer directly corresponds to the number of concepts that can be explicitly represented and manipulated.
@@
$\mathcal{L}_{\text{{SAEmnesia}}}=\mathcal{L}_{\text{unsupSAE}}+\beta\mathcal{L}_{\text{supSAE}}+\lambda\mathcal{L}_{\text{L1}}.$ (6)
@@
Where $\mathcal{L}_{\text{unsupSAE}}$ is the loss as defined in Eq. 3, $\mathcal{L}_{\text{supSAE}}$ is the supervised loss that enforces the desired concept-latent relationships and $\mathcal{L}_{\text{L1}}$ is the sparsity regularization term. $\mathcal{L}_{\text{supSAE}}$ consist of our Concept Assignment (CA) loss with a weighted additional Decorrelation (DC) loss:
$\mathcal{L}_{\text{supSAE}}=\mathcal{L}_{\text{CA}}+\gamma\mathcal{L}_{\text{DC}}.$ (7)
@@
Concept assignment loss. Unlike traditional approaches that only apply reconstruction loss globally across all latents, our method strengthen concept-latent bonds that enable one-to-one mappings. Our CA loss is computed exclusively for latents that are assigned to concepts present in each training sample, encouraging these specific latents to activate strongly. Given a training sample, we define the binary ground-truth vector $\mathbf{y}=[y_{1},\ldots,y_{K}]^{\top}\in\{0,1\}^{K}$, where $y_{k}=1$ if concept $c_{k}$ is present in the sample, and 0 otherwise. We denote with $\mathcal{T}$ the set of concepts present in that sample. The CA loss measures how much the assigned latents activate when their corresponding concepts are present:
$\mathcal{L}_{\text{CA}}=\frac{1}{B}\sum_{b=1}^{B}\frac{1}{|\mathcal{T}^{(b)}|}\sum_{c\in\mathcal{T}^{(b)}}\big[-\log\sigma(v^{(b)}_{i_{c}})\big$], (8)
where $v_{i_{c}}$ represents the pre-activation (logit), and $\sigma(\cdot)$ is the sigmoid function.
Decorrelation constraint. To promote disentanglement across multiple macro-categories of concepts, we generalize the decorrelation constraint to $M$ disjoint concept groups. Specifically, we partition the full concept set into non-overlapping subsets $\mathcal{C}=\bigcup_{m=1}^{M}\mathcal{C}_{m}$, with $\mathcal{C}_{m}\cap\mathcal{C}_{m^{\prime}}=\varnothing$ for $m\neq m^{\prime}$, where each $\mathcal{C}_{m}$ represents a high-level group of related concepts (e.g., objects, styles, materials, or other semantic categories depending on the dataset). We denote by $\mathcal{I}_{\mathcal{C}_{m}}=\{\Phi(c)\mid c\in\mathcal{C}_{m}\}$ the set of latent indices assigned to the concepts in group $\mathcal{C}_{m}$. Given a mini-batch of size $B$, we define the activation vector of each concept $c$ as $\mathbf{a}_{c}=[v^{(1)}_{i_{c}},v^{(2)}_{i_{c}},\ldots,v^{(B)}_{i_{c}}$$^{\top}$. The multi-group decorrelation constraint is then formulated as:
$\mathcal{L}_{\text{DC}}=\frac{\displaystyle\sum_{m<m^{\prime}}\sum_{i\in\mathcal{I}_{\mathcal{C}_{m}}}\sum_{j\in\mathcal{I}_{\mathcal{C}_{m^{\prime}}}}\rho(\mathbf{a}_{i},\mathbf{a}_{j})}{\displaystyle\sum_{m<m^{\prime}}|\mathcal{I}_{\mathcal{C}_{m}}|\,|\mathcal{I}_{\mathcal{C}_{m^{\prime}}}|}.$ (9)
Here, $\rho(\mathbf{a}_{i},\mathbf{a}_{j})$ denotes the Pearson correlation coefficient between activation vectors $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$. This constraint penalizes correlations between activation patterns of latents assigned to different concept groups.
@@
Sparsity regularization. To encourage sparse activations in the latent representation, we incorporate an L1 regularization term that penalizes the magnitude of latent activations $\mathbf{v}$. This sparsity constraint promotes the emergence of interpretable features by encouraging most latents to remain inactive for any given input, thereby improving the disentanglement of learned representations.